OpenStudy (anonymous):

Solve the following system. y = 3x^2 + 5x -1 y = 4x + 3

OpenStudy (anonymous):

Use substitution. y = y, so: 3x^2 + 5x -1 = 4x +3 (move everything to the left) 3x^2 +x - 4 = 0 (now factor) (3x + 4)(x - 1) = 0 x = -4/3 x = 1